45 candidates appear in a competitive examination. Prove that there are at least two candidates

whose roll numbers differ by a multiple of 44.


1
Expert's answer
2021-11-26T05:15:36-0500

Taking into account that there are 44 different remainders of the division by 44, we conclude by Pigeonhole Principle that among 45 roll numbers thare are at least two numbers "a" and "b" with the same remanders. It follows that "a=44t+r" and "b=44s+r" for some "t,s\in\Z" and "0\le r<44." Therefore, "a-b=(44t+r)-(44s+r)=44(t-s)." We conclude that 44 divides "a-b," and hence there are at least two candidates (with roll numbers "a" and "b") whose roll numbers differ by a multiple of 44.